In a triangle, one angle measures 44 degrees. the measure of the second angle is twice the measure of the third angle. Find the measure of the third angle
Found 2 solutions by macston, Boreal:
Answer by macston(5194)

.
F=first angle=44 degrees; T=third angle; S=second angle=2T
.
F+S+T=180 degrees
44 degrees + 2T + T=180 degrees
3T=136 degrees
T=45.33 degrees
ANSWER: The third angle measures 45.33 degrees.
.
CHECK:
S=2T=2(45.33)=90.66 degrees
F+S+T=180 degrees
44 degrees + 90.66 degrees + 45.33 degrees=180 degrees
180 degrees=180 degrees

Answer by Boreal(15235)

Sum is 180 and one is 44, so the other two are 180-44=136 degrees.
x=smaller angle
2x=larger
2x+x=136
3x=136
x=136/3=45 (1/3) degrees. Third angle
(second angle is 90 2/3 degrees, Those two sum to 136 degrees.